OpenStudy (anonymous):

I need to solve this equation for t: 0.002=0.1+e^(t/2000), which after subtracting 0.1 from both sides I am left with -0.098=e^(t/2000). I don't have a clue how to solve it. Do I need to take the Ln of both sides?

OpenStudy (kirbykirby):

yes

OpenStudy (kirbykirby):

ah wait you'll get a negative argument for the ln

OpenStudy (kirbykirby):

It does appear like you'll get some complex logarithm

OpenStudy (anonymous):

maybe I have something wrong, this is the second part of a problem, which was right for the first question.

OpenStudy (kirbykirby):

Did you every use complex numbers to solve these? Or maybe you did do a wrong calculation if you expect a real (non-complex) answer

OpenStudy (kirbykirby):

ever*

OpenStudy (anonymous):

No I don't think I should end up with complex numbers, I'm looking for an exact time in minutes.

OpenStudy (anonymous):

The problem goes like this: A swimming pool whose volume is 10,000 gal contains water that is 0.001% chlorine. Starting at t=0, city water containing 0.001% chlorine is pumped into the pool at a rate of 5 gal/min. The pool water flows out at the same rate. What is the percentage of chlorine in the pool after 1 hour? When will the pool water be 0.002% chlorine?

OpenStudy (kirbykirby):

is the % of chlorine initially and being pumped the same?

OpenStudy (anonymous):

yes, the input and output are the same rate. I was able to solve the percentage of the chlorine in the pool after 1 hour was I got the general solution x(t)=0.1+0.9/e^(t/2000) and then using that to solve for x(60min). which I got 0.0097% which was correct,

OpenStudy (kirbykirby):

SOrry I meant the percentage %, not the rate of flow

OpenStudy (anonymous):

yes the pool contains 0.01% chlorine initially. Is that what your asking?

OpenStudy (kirbykirby):

ah ok.

OpenStudy (kirbykirby):

What's the 2nd part of the problem then

OpenStudy (anonymous):

so the second part of the question that I am having trouble with is, When will the pool water be 0.002% chlorine? which I just plugged 0.002% in for x(t) and got 0.002=0.1+0.9/e^(t/2000). But now I'm wondering if that is right

OpenStudy (kirbykirby):

Hm I get: \(\Large 0.1+0.9e^{-t/2000}=0.973 gallon\) but then \[ \Large \frac{0.973}{10,000}=0.0000973=>0.00973\%\] So similarly, for 0.002%, it should give 0.2 gallon to plug into the equation.

OpenStudy (anonymous):

interesting, I just lucked out because I just divided 0.973 by 100. So my solution is wrong. But why, or how did you multiply e^(t/2000) to C?

OpenStudy (kirbykirby):

Not sure what you mean? Did you get a step like: \[\ln|0.1-x|=\frac{-t}{2000}+C\]

OpenStudy (anonymous):

I was referring to the general solution, When I solved for (mu) and then used mu to get the general equation. \[d/dt[(x)e^(t/2000)]=0.00005e^(t/2000)\]

OpenStudy (anonymous):

because I integrated both sides which gave me xe(t/2000)=0.1e^(t/2000)+C, so I divided both sides by e^(t/2000) which is how I got x=0.1+C/e^(t/2000).

OpenStudy (kirbykirby):

Hm, I did the following way: \[ \frac{dx}{dt}=0.00005-5\frac{x}{10000}\\ =0.00005-0.0005x\\ =0.0005(0.1-x)\\ \frac{1}{0.1-x}\frac{dx}{dt}=0.0005\\ \frac{1}{0.1-x}dx=\frac{1}{2000}\,dt\\ -\ln{|0.1-x|}=\frac{t}{2000}\\ \ln{|0.1-x|}=\frac{-t}{2000}+C\\ |0.1-x|=e^{-t/2000}e^C\\ 0.1-x=\pm e^Ce^{-t/2000}=Ae^{-t/2000}\\ x=0.1-Ae^{-t/2000}\] Now at the intial condition, \(x(0)=1\), since 0.01% into gallon, just like before, is 1 so: \(x(0)=1=0.1-Ae^0\implies A=-0.9\) , soo: \[ x=0.1+0.9e^{-t/2000}\]

OpenStudy (anonymous):

wow, that is much easier than what I was trying to do. I was trying to set it up the way my professor showed us by using the standard form, getting p(t), and Q9t), and mu(t). Thank you so much.

OpenStudy (kirbykirby):

:) you're welcome.

OpenStudy (kirbykirby):

You shouldn't get a negative argument for the ln now :) especially since you have to be careful about the percentage thing .

OpenStudy (anonymous):

right, thanks again.